Solve each system by substitution. Show your work.
1) a= 3b+1
5a-3b= 13
2) 4y= -3x
5x+8y= 4
3) y= 2x
x+y= 21
4) r= -3s
3r+ 4s= -10
5) y= x+3
3x+ 2y= 26
6) a+b= 11
3a-2b= 8
7) 7x-3y= 23
x+2y= 13
8) 2x+3y= 7
4x-5y= 25

Answer by Edwin McCravy(20086)

I'll just do 5 and 8 for you:

5) y= x+3
 3x+ 2y= 26 

Copy the second equation, but instead of y, don't 
write y, but instead write what y equals from the
first equation, which is x+3 in parentheses (x+3)

 3x+ 2(x+3)= 26

Remove parentheses by distributing

 3x + 2x + 6 = 26
      5x + 6 = 26
          5x = 20
           x = 4

Substitute (4) for x in y= x+3

           y= 4+3
           y = 7

Solution: (x,y) = (4,7)

----------------------------

8) 2x+3y= 7
 4x-5y= 25

We notice that the second equation has 4x and the first
equation has 2x, and we know that 4x is 2 times 2x

So we solve the first equation for 2x

   2x+3y= 7

We subtract 3y from both sides

      2x= 7-3y

We need 4x, not 2x, to substitute in the second equation,
so we multiply both sides, every term, of that equation by 2:

      4x= 14-6y

Copy the second equation,  4x-5y= 25 but instead of 4x, 
don't write 4x, but instead write what 4x equals from that
equation, which is 14-6y.

      14-6y-5y= 25
Combine like terms
        14-11y = 25 

Add -14 to both sides:

          -11y = 11
             y = -1

Solution: (x,y) = (5,-1)

Edwin